The first term in the arithmetic sequence is 1 and the fifth is 3. What is the tenth term in the sequence?


1
Expert's answer
2021-12-20T16:49:36-0500
"a=1"

"a_5=a+4d=3"

"1+4d=3"

"d=0.5"

"a_{10}=a+9d=1+9(0.5)=5.5"

"a_{10}=5.5"